Explain in words how to write an equation that is part one: parallel and then also part two: perpendicular to the equation y=(2/3)x-4 passing through the point (-2,-5). Write your answer in standard form.
Answer by TimothyLamb(4379)

---
parallel lines have equal slopes:
m = 2/3
point(-2,-5)
---
point-slope form:
y + 5 = 2/3(x + 2)
standard form:
y + 5 = (2/3)x + 4/3
(2/3)x - y = -4/3 + 5
(2/3)x - y = -4/3 + 15/3
answer:
(2/3)x - y = 11/3
---
---
perpendicular lines have negative reciprocal slopes:
m = -3/2
point(-2,-5)
---
point-slope form:
y + 5 = -3/2(x + 2)
standard form:
y + 5 = (-3/2)x - 3
(-3/2)x - y = 3 + 5
answer:
(-3/2)x - y = 8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php